OpenStudy (anonymous):

i am working this problem: (2^30-2^29)/ 2. The book says the answer is simplified to 2^30=2^30 how?

OpenStudy (amistre64):

well, i spose it could also simplify to 15=15 , or 1/2 = 1/2, or ...

OpenStudy (amistre64):

we can factor out a 2^29 from the top ...

OpenStudy (anonymous):

i forgot part of the problem: (2^30-2^29)/2 =2^28 The books says the answer is simplified to 2^30=2^30 . I get 2^29=2^29

OpenStudy (anonymous):

(2^30-2^29)/2 =2^28 This is correct.